OpenStudy (anonymous):

cos^2(θ)(1+tan^2(θ))=1 I need to prove this: so far I got that tanx=sinx/cosx: cos^2(θ)(1+sin^2(θ)/cos^2(θ)) But I am not sure if that is correct

OpenStudy (anonymous):

= ( 1/ sec²θ ) * sec²θ = 1

OpenStudy (amistre64):

1+tan^2 = sec^2 = 1/cos^2

OpenStudy (amistre64):

sin^2 + cos^2 = 1 sin^2 + cos^2 = 1 ---------------- = tan^2 + 1 = sec^2 cos^2

OpenStudy (anonymous):

you can also multiply the left side out to get cos^2x + sin^2x.... = 1